Shift summary for evening shift, July 10, 99

B. Cox (cox@fnal.gov)
Sun, 11 Jul 1999 00:46:57 -0500 (CDT)

Shift Summary for Evening Shift, 10/Jul/99

On Shift: B. Cox, Hogan

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 8.0
Hours of E832 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

TeV LCW Leak RF Beam Off 15:30 - 16:07
Changing to pos pol./start run 16:15 - 16:25

Changing tapes/start run 13853 22:00 - 22:20

Bad CsI channel 2575 23:35 - 00:15
Access to replace DPMT
(2231 replaced by 3260)
Hogan/Brad

========================================================================
Runs ended this shift:

Run 13851 Muon Run
Run 13852 E832_Beam
Run 13853 E832_Beam

Tapes taken to FCC this shift

QKK110-112
QKS110-112
QKC110-112
QKB038

========================================================================
Accesses during shift:

One by Hogan/BC to rep[lace DPMT 2575

========================================================================
Overview of shift:

17:10 CAMAC error indication from SDAQ ; Crate 23 DTM front panel
is on "tracking".

17:12 Dave filled alcohol bubbler

17:30 Ed calls. Channel 1247 is drifting. He wants to
be called if it gets beyond 5%.

17:40 ktevana error; ran kfrend -> success

18:15 removed split tape QKNL08; mounted QKNL09

20:13 Two red buttons on SDAQ

CAMAC error indication from SDAQ ;
Crate 10 DTM front panel is on "tracking".

Epicure NM1U magnet current slightly out of spec.

Both clear.

21:10 Request for mounting QKSM05. Done.

21:50 ktevana error - ran kfrend

22:00 Tapes full - change tapes

22:20 Run started

23:35 Spill 54 - channel 2575 has pedexp>0.
resetting the pipeline several times w/DBC does not help.
stop run to swap the DPMT

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13851 16:07 - 16:15 Muon Run
13852 16:26 - 22:00 E832_Beam Magnet Positive
13853 22:20 - 24:25 E832_Beam Magnet Positive
Stopped run because of DPMT 2575
problem
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

A few cosmics counters have wide pedestal RMS's, causing them to be readout
more often. Valeri and Val notified.